How many 3 number combinations can be made with these 6 numbers 242200.
Answer by Edwin McCravy(20081)

Since order matters, we should not use the word "combination", In
algebra, the word  "combination" is used only when the order does 
not matter.

But the order does matter here, so we should use the word "permutation"
instead of "combination"

Case 1:  None of the digits are the same:

These are the arrangements of {0,2,4}

Choose the first digit 3 ways
Choose the second digit 2 ways
Choose the third digit 1 way.

3×2×1 = 6 ways for case 1.  Those 6 are

024, 042, 204, 240, 402, 420
--------------------------------------------
Case 2:  Exactly two 0's

Choose the remaining digit 2 ways.  {2,4}

We can arrange them in 3!/2! = 3 ways

That's 2×3 = 6 for case 2.  Those 6 are:

002, 004, 020, 040, 200, 400
-------------------------------------------
Case 3.  Exactly 2 2's:

Choose the remaining digit 2 ways.  {0,4}

We can arrange them in 3!/2! = 3 ways

That's 2×3 = 6 for case 3.  Those 6 are:

022, 224, 202, 242, 220, 422
------------------------------------------
Case 4.  All 3 2's:

That's just 1 way, 222.
------------------------------------------
Grand total = 6+6+6+1 = 19 ways.

Edwin